A spark of 20MV lightning has an energy output of 0.150 mw-s. Compute how much charge flowed through the lightning.


1
Expert's answer
2021-04-23T11:07:12-0400
"q=\frac{W}{U}\\q=\frac{0.15\cdot10^6}{20\cdot10^6}=0.0075\ C"